The soil in a certain area has on average 2.9 mg/mA3 of substance A. If 5 soil tests show a concentration of a substance A of 1.8, 2.3, 2.9, 3.4 and 4.1 mg/m^3 in the soil, construct a 95% confidence interval for ơ2. Assume that the population is normally distributed

need help with Statistics assignment
Answer #1

ar x = (1.8 + 2.3 + 2.9 + 3.4 + 4.1)/5 = 2.9

s2 = ((1.8 - 2.9)^2 + (2.3 - 2.9)^2 + (2.9 - 2.9)^2 + (3.4 - 2.9)^2 + (4.1 - 2.9)^2)/4 = 0.815

At 95% confidence interval the critical values are a/2,n-1 = X0.025.4 = 11.1433

                                                                              1-a/2,n-1 = 0.075.4 = 0.4844

The 95% confidence interval for sigma^2 is

(n - 1)s2/a/2,n-1 < sigma^2 < (n - 1)s2/1-a/2,n-1

= 4 * 0.815/11.1433 < sigma^2 < 4 * 0.815/0.4844

= 0.29 < sigma^2 < 6.73